# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v1

@@ 7 Discussions
To further validate the robustness and versatility of our proposed framework, we conduct three systematic analyses addressing critical operational aspects of crop mapping tasks: temporal adaptability, and data efficiency. These experiments are designed as follows.
  • Quantify the model’s capacity to handle variable-length temporal sequences—a fundamental requirement for crop mapping across diverse phenological cycles and cloud cover conditions. We systematically evaluate its performance on early-season crop mapping using temporally downsampled input sequences. Specifically, we subsample the original Sentinel-2 time series into fixed-interval sequences of 8-, 12-, and 16-step configurations. This simulates scenarios with varying observational frequencies while preserving phenological relevance.
@@
The label efficiency analysis (Table 9 and Figure 13 c)) underscores the practical value of our foundation model pretraining strategy. As the fine-tuning dataset scales from 10% to 100%, the pretrained variant demonstrates consistent performance gains, achieving a 57.63% F1 score with only 50% fine-tuning data—surpassing the non-pretrained model’s full-data performance (54.82%). This highlights the pretraining framework’s ability to extract transferable spatiotemporal representations, enabling effective adaptation to downstream tasks with limited annotations. Notably, while the non-pretrained model plateaus beyond 50% fine-tuning data ($\Delta$F1=+0.76from 50% to 100%), the pretrained variant continues to improve ($\Delta$F1=+2.25%), suggesting superior utilization of additional labeled samples.
@@ 8 Conclusion
This paper first addresses and validates the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping. It leads to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial scaling operations. Capitalizing on this finding, we introduce AgriFM, a multi-source remote sensing foundation model that adapts Video Swin Transformer with specialized spatiotemporal modeling for agricultural crop mapping. AgriFM integrates multi-source data including MODIS, Landsat-8/9, and Sentinel-2 temporal satellite data, and employs geographical prior land cover products for supervised pre-training. It effectively addresses the limitations of existing RSFMs, ensuring the incorporation of multi-resolution data, providing comprehensive coverage of the crop growth cycle, optimizing the use of geographical information, and offering a unified framework for a wide array of crop mapping tasks. These tasks include, but are not limited to, cropland mapping, cropland boundary delineation, early-season crop mapping, and specific tasks like wheat and rice mapping. The effectiveness and adaptability of AgriFM have been substantiated through superior performance across a variety of crop mapping tasks, demonstrating its potential as an immediately applicable solution for leveraging deep learning in more precise and efficient crop mapping.
@@
Despite the agricultural applications, the natural extension of this work lies in assessing the model’s generalizability to other Earth observation tasks. Many geo-spatial applications similarly require simultaneous processing of spatiotemporal data, suggesting our foundation model could serve as a versatile backbone. However, task-specific adaptations—particularly in decoder design—remain necessary. For vegetation parameter estimation (e.g., LAI and FVC), for instance, the decoder must generate time-specific outputs rather than integrated predictions. While Video Swin Transformer outperforms other architectures for crop mapping, performance variations across tasks are substantial. For certain image classification or regression tasks, ViT-based methods may still prevail due to their global representation capabilities.
@@
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 7 Discussions
To further validate the robustness and versatility of our proposed framework, we conduct more systematic analyses addressing critical operational aspects of crop mapping tasks: temporal adaptability, and data efficiency. These experiments are designed as follows.
  • Quantify the model’s capacity to handle variable-length temporal sequences—a fundamental requirement for crop mapping across diverse phenological cycles and cloud cover conditions. We systematically evaluate its performance on early-season crop mapping using temporally downsampled input sequences. Specifically, we down-sample the original Sentinel-2 time series into fixed-interval sequences of 8-, 12-, and 16-step configurations. This simulates scenarios with varying observational frequencies while preserving phenological relevance.
@@
The label efficiency analysis (Table 9 and Figure 13 b)) underscores the practical value of our foundation model pretraining strategy. As the fine-tuning dataset scales from 10% to 100%, the pretrained variant demonstrates consistent performance gains, achieving a 57.63% F1 score with only 50% fine-tuning data—surpassing the non-pretrained model’s full-data performance (54.82%). This highlights the pretraining framework’s ability to extract transferable spatiotemporal representations, enabling effective adaptation to downstream tasks with limited annotations. Notably, while the non-pretrained model plateaus beyond 50% fine-tuning data ($\Delta$F1=+0.76from 50% to 100%), the pretrained variant continues to improve ($\Delta$F1=+2.25%), suggesting superior utilization of additional labeled samples.
@@ 8 Conclusion
This paper first addresses and validates the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping. It leads to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial down-sampling operations. Capitalizing on this finding, we introduce AgriFM, a multi-source temporal remote sensing foundation model with specialized spatiotemporal modeling for agricultural crop mapping. AgriFM integrates multi-source satellite data including MODIS, Landsat-8/9, and Sentinel-2, and employs geographical prior land cover products for supervised pre-training. It effectively addresses the limitations of existing RSFMs, ensuring the incorporation of multi-resolution data, providing comprehensive coverage of the crop growth cycle, optimizing the use of geographical information, and offering a unified framework for a wide array of crop mapping tasks. These tasks include, but are not limited to, cropland mapping, cropland boundary delineation, early-season crop mapping, and specific tasks like winter wheat and rice mapping. The effectiveness and adaptability of AgriFM have been substantiated through superior performance across a variety of crop mapping tasks, demonstrating its potential as an immediately applicable solution for leveraging deep learning in more precise and efficient crop mapping.
@@
Despite the agricultural applications, the natural extension of this work lies in assessing the model’s generalizability to other Earth observation tasks. Many geo-spatial applications similarly require simultaneous processing of spatiotemporal data, suggesting our foundation model could serve as a versatile backbone. However, task-specific adaptations—particularly in decoder design—remain necessary. For vegetation parameter estimation (e.g., LAI and FVC), for instance, the decoder must generate time-specific outputs rather than integrated predictions. While Swin Transformer outperforms other architectures for crop mapping, performance variations across tasks are substantial. For certain image classification or regression tasks, ViT-based methods may still prevail due to their global representation capabilities.